6x-7=2x+29 x=9



How is this?

Found 2 solutions by CassieLynn, stanbon:
Answer by CassieLynn(3)

6(9)-7=2(9)+29
54-7=18+29
47=47

Answer by stanbon(75887)

Too many equality signs.
Cheers,
Stan H.